find the value of log3log3log9log729+3log3 basev3

Grade:11

1 Answers

Sher Mohammad IIT Delhi
10 years ago
9^3=729
log9(729)=3 3*logv3((v3)^2)=3*2=6
log3(3)=1
log3(1)=0
log3log3log9729+3logv33=6
Sher Mohammd
B.Tech IIT Delhi